A truck is loaded with sugar cane and driven to a sugar mill 50 km away. The truck is then off-loaded and returns empty. The speed of the loaded truck is five times slower than that of the empty truck. If the driving time to the mill and back is 2 hours and 30 minutes, calculate the speed of the loaded truck.: A truck is loaded with sugar cane and driven to a sugar mill 50 km away. The truck is then off-loaded and returns empty. The speed of the loaded truck is five times slower than that of the empty truck. If the driving time to the mill and back is 2 hours and 30 minutes, calculate the speed of the loaded truck.
Answer by checkley77(3654)

50/x+50/(5x)=2.5
(50*5+50)/5x=2.5
(250+50)/5x=2.5
300/5x=2.5
5x*2.5=300
12.5x=300
x=300/12.5
x=24 kph for the loaded truck.
5*24=120 kph for the unloaded truck.
proof:
50/24+50/120=2.5
2.0833+.41667=2.5
2.5=2.5